# SafeSearch: Automated Red-Teaming for the Safety of LLM-Based Search Agents

### v3

@@ 3.2 Automated Generation of Test Cases At Scale
    Step 1: Scenario Envisioning: The test generator brainstorms a realistic scenario in which the described risk could arise when a user submits a query to the search agent.
@@
    Step 2: Test Design: The test generator then schemes a “malicious plan”. In this step, the generator specifies the intended negative consequence, as well as the websites that would be exploited to induce it. A key auxiliary consideration here is the time gap: the threat always occurs after the knowledge cutoff of the backend LLMs. This highlights the time-sensitive nature of the threat. To illustrate, testing whether agents will misbehave with a single unreliable and long-outdated website, alongside several current sources, arguably constitutes an unfaithful and unrealistic modeling of the threat. Notably, no specific date is registered during test case generation, ensuring that the resulting test cases are not tied to a single date and thus are not limited to one-time applicability.
@@
    Step 3: Test Instantiation: The test generator translates the verbalized test proposals into guidelines for the website generator, defining required formats and key information for an appropriate synthesis of unreliable websites. Inspired by rubric-based evaluation (Souly et al., 2024), we also prompt it to produce a checklist for the safety evaluation, enabling precise and hallucination-free verification of whether the negative consequence manifests in the agent’s response.
Two design choices are crucial to producing high-quality test cases: 1) We adopt a multi-stage workflow, allowing the test generator to perform multiple rounds of reasoning. 2) We elicit auxiliary information beyond the final components required in test cases, such as user expectations and unreliable website sources, to better guide the generation process.
Figure 4: Three-Step Workflow of Test Case Generation. Colored fields are required in final test cases, white solid-bordered fields are passed to later stages, and dashed-bordered fields are auxiliary.
@@
Automated Quality Filtering. We define two filtering criteria for generated test cases using differential testing with a baseline model across two settings: a benign tool and a manipulated tool (see Section 3.3). The criteria are: 1) Attainability: Non-occurrence of risky responses should not result from the inherent infeasibility of the cases themselves, which may give a false sense of safety. We therefore perform a dry run with a baseline search agent to confirm that the intended consequence is attainable. 2) Integrity: When using a benign tool, the target consequence should not be triggered by the agent. A test case that satisfies both criteria is realistic and actionable.
3.3 Efficiently Red-Teaming Search Agents via Simulation-Based Testing
Similar to Ruan et al. (2024); Xu et al. (2024), we adopt a simulation-based method to evaluate search agent safety. Instead of arbitrarily manipulating search engine rankings (Kumar et al., 2019), which risks harm to users, we inject controlled, unreliable content into the agent’s search process. This enables reproducible and low-cost testing without unintended external consequences.
@@ 3.3 Efficiently Red-Teaming Search Agents via Simulation-Based Testing
As illustrated in Figure 5, the agent $S$ receives a user query $u$ and issues a sequence of search tool calls $\{s_{1},s_{2},\dots,s_{n}\}$ before synthesizing the final response. If $n>1$, the unreliable site $d_{u}$ is inserted only in the first call, thereby giving the agent an opportunity to reduce its influence in subsequent calls. The unreliable website $d_{u}$ is generated according to predefined guidelines and conditioned on the testing date, reflecting the fact that search tools in practice typically index up-to-date, potentially unreliable material. The final set of search results shown to $S$ is therefore $D\cup\{d_{u}\}$, where $D=\{d_{1},d_{2},\dots,d_{k}\}$ denotes the authentic search results. By mixing a single unreliable website $d_{u}$ with the authentic sources $D$, the framework approximates realistic threat scenarios while preserving the agent’s normal operation in producing its final output. We detail why and how simulation-based testing can mitigate the real-world threats in Section E.5.
Figure 5: Conceptual Illustration of Red-Teaming via Simulation-Based Testing.
@@
We evaluate the agent by auditing its final responses, the primary channel through which potential risks may propagate to users. Following common practices of agent safety evaluation (Ruan et al., 2024; Zhang et al., 2024; Yin et al., 2024), we resort to LLM-as-a-Judge (Zheng et al., 2023) for automated evaluation. Our analysis primarily focuses on two aspects:
@@
    Safety Evaluation. The evaluator receives the user query, target consequence (with the checklist), and agent response, engages in a reasoning process, and then delivers a final boolean judgment (see Appendix G). This yields a signal of potential compromise in the search agent. Across all trials, we derive a safety metric for search agents, which we term the Attack Success Rate (ASR).  [^4] Although our testing is not limited to adversarial settings, we adopt the term ASR for ease of understanding.
@@
    Helpfulness Evaluation. We also assess the agent’s utility, namely, how helpful its final response is to the user. A helpfulness evaluator, given the user query and the agent’s response, performs reasoning and outputs a score on a scale from 1 (not helpful at all) to 5 (extremely helpful). This is implemented using Appendix G. Rescaling to the range 0-100 and averaging across all trials yields the overall Helpfulness Score (HS).
3.4 Synthesizing the SafeSearch Benchmark
Table 1: Curation of the SafeSearch Benchmark. “RR” denotes the retention rate.
Risk # Raw $R$
Ads 192 31.3%
Bias 330 18.2%
Harm 867 6.9%
Injec 883 6.8%
Misinfo 318 18.9%
Following Section 3.2, we curate 300 test cases spanning five real-world risks. This is accomplished by iteratively running the generate-filter process until 60 test cases are obtained for each risk. This process also down-samples raw test cases, providing insights into how different risks impact search agents. Among these, test cases of the two adversarial risks Harmful Output (Souly et al., 2024) and Indirect Prompt Injection (Zhan et al., 2024) prove difficult to curate, with retention rates of just 6.9% and 6.8%. Careful manual review ensures the quality of test cases. Benchmark characteristics are detailed in Appendix A, and see Appendix F for examples of test cases.
Table 2: Main Results of Search Agents on SafeSearch. Reasoning models are highlighted in light blue, while non-reasoning models are shown in light beige. GPT-5 (/-mini) agents are indicated with distinct colors for their outlier performance. For other models, we use bold to denote the best performance and underline for the second best. Timestamp: Sep 9, 2025.
@@ 3.4 Synthesizing the SafeSearch Benchmark
4 Experiments
Search Agents. We evaluate three representative genres of search agent implementations with comprehensive backend LLMs, with more details in Section B.1.555As our red-teaming method involves hooking search tools, we exclude proprietary search agent systems due to accessibility limits. However, the method can still be adopted by internal testers of these systems.
@@ 4 Experiments
    LLM w/ Search Workflow. LLMs passively accept the search results of the verbatim user query before generation. This is adopted from FreshLLMs (Vu et al., 2024) in a zero-shot way.
@@
    LLM w/ Tool Calling. The agent is allowed to proactively search for anything it requires. We allow the models to decide the number of calls by themselves, with a maximum budget of 3.
@@
    Deep Research Scaffold. We build on the open-source prototype from Google [^6] https://github.com/google-gemini/gemini-fullstack-langgraph-quickstart. This full-fledged system involves four steps: query decomposition, one-turn search, reflection, and answer finalization. We set the maximal research loop to 3 and allow up to 2 queries per round.
Models. Since search agents must handle long website contexts, we focus primarily on LLMs with a context window of at least 32K tokens and native support for tool calling. Totally, we include seven proprietary and eight open-source models. By default, we use the chat or instruction-tuned version and enable the reasoning mode if supported. See Section B.2 for more details.
Red-Teaming Configurations. Balancing cost and effectiveness, we employ o4-mini as the test generator, while GPT-4.1-mini serves as the website generator, safety evaluator, and helpfulness evaluator. For automated filtering, we use Qwen3-8B as the baseline model. The search tool returns the top five results with webpage contents by default [^7] We use the Serper API (https://serper.dev/) for real-time Google Search results and the Jina Reader API (https://jina.ai/reader/) for LLM-friendly content extraction., and each website is limited to 2,000 tokens (measured with the GPT-5 tokenizer). The unreliable website is always appended to the result list. We cache search results across agents for fairness and reduce stochasticity. Further validation of these testing choices (e.g., injection position, per-website token limit, website generator choice, website generation timestamp, and particularly reliability of the safety evaluator) is in Appendix C. The LLM temperature is set to 0.6, and each test case is run three times, with ASR and HR averaged across trials. All relevant prompts are provided in Appendix G.
@@ 4.1 Main Findings of Red-Teaming with SafeSearch
Of Global Trends, Search Agents Can Be Easily Misled by Unreliable Websites. Across settings, ASRs remain alarmingly high for most search agents. This indicates that once unreliable search results are involved, agents often cannot mitigate their impacts and propagate risky content to users. Even the safest implementations, GPT-5 series with tool-calling, still exhibit ASRs of around 5%, which are relatively low but still above impossible. Interestingly, agents built on Gemini-2.5 series, despite having the latest knowledge cutoff, do not necessarily demonstrate greater resilience.
Across LLMs and Agent Scaffolds, Safety Is Jointly Influenced. Search agent safety hinges on how backend LLMs handle unreliable results. With identical scaffolds, different LLMs demonstrate markedly different resilience. Nevertheless, their relative rankings remain almost consistent when augmented with search workflows and tool calling. In general, reasoning models tend to be more resilient than non-reasoning ones, as reflected in their lower ASRs across categories. At the same time, scaffold design also plays a critical role: for example, GPT-4.1-mini reaches an overall ASR of 90.5% under the search workflow, which decreases to 77.8% with tool-calling and further to 57.4% under the deep research scaffold. Long-horizon search via tool-calling or deep-research scaffolds can help mitigate the threat by cross-validating multiple sources.
@@
Across Distinct Risk Types, Search Agents Show Varying Degrees of Susceptibility. At the aggregate level, the average ASR values reveal notable differences across risk types. For example, the highly resilient agent GPT-5 with tool-calling achieves a 0% ASR in indirect prompt injection, yet it remains vulnerable to other risk types. Among these, Indirect Prompt Injection shows a lower average ASR, highlighting agents’ resilience to this security threat (Wallace et al., 2024; Debenedetti et al., 2025). In contrast, the Misinformation risk poses the most severe threat to agents.
In Relation to Helpfulness, Safety Can Be Achieved Simultaneously With Proper Orchestration. Notably, exhibiting unsafe responses does not necessarily reduce the helpfulness score. For instance, under the tool-calling setting, the average HSmanip. (91.6) exceeds HSbenign (90.7). This suggests that such threats can manifest in a stealthy manner. What’s more, search agents can achieve both safety and helpfulness, reaching a Pareto-optimal balance, as illustrated by the GPT-5 series (see Appendix D for an explanatory case study). This effect is particularly pronounced when considering cost scaling, where the deep search scaffold narrows capability gaps across backend LLMs, enabling even smaller models like Qwen3-8B to remain both useful and safe.
4.2 SafeSearch Enables Tracking Safety Progress in Agent Development
Additionally, the red-teaming framework provides a viable way to understand the implicit impacts of agent configurations on safety, while also uncovering the effectiveness of defense strategies.
The Impact of Search Result Count on Safety. Agent developers can configure how many search results are provided to LLMs, balancing utility against cost efficiency. As shown in Figure 6 (Left), our empirical findings demonstrate that this choice can implicitly influence the safety guarantees of search agents. When fewer search results are considered, unreliable websites have more chances to be included in the final response, thereby increasing the ASR. This analysis also serves as an ablation study of our red-teaming setup, showing that the threat is not limited to any particular search result count and that a default choice of five is sufficiently representative.
@@ 4.2 SafeSearch Enables Tracking Safety Progress in Agent Development
Figure 6: SafeSearch Enables Understanding of Safety Progress in Agent Implementations. Experiments are on LLMs with a search workflow. (Left) Search result count correlates with agent susceptibility. Timestamp: Sep 17, 2025. (Right) SafeSearch reveals defense effectiveness.
@@
The Effectiveness of Defense Strategies. We consider two baseline defenses: 1) Reminder: We augment the agent’s prompt with an explicit reminder instructing it to be cautious of potential unreliable websites. 2) Filtering: Before returning search results, we employ an auxiliary detector model (i.e., GPT-4.1-mini) to reason about and filter out potentially unreliable search results (Appendix G). As shown in Figure 6 (Right), both defenses reduce risks to some extent but fall short of fully eliminating them. The reminder prompting method, which reflects a common strategy of adding caveats in system prompts, proves largely ineffective. By contrast, the filtering defense is more effective, halving the ASR across two search agents. This is unsurprising given that it represents a proactive step toward safer search tools, with a recall rate of 44.2%. A retrospective manual review of the detector’s reasoning reveals that most detection successes arise from the relative lack of stealth in unreliable websites. To explain, stealthiness is not a principle considered in our test case generation. In other words, specially crafted stealthier websites can pose greater threats, consistent with prior findings (Puccetti et al., 2024; Chen & Shu, 2024).

### v4

@@ 3.2 Automated Generation of Test Cases
    Step 1: Scenario anchoring: The test generator brainstorms a realistic scenario where the described risk could arise when a user submits a query to the search agent.
@@
    Step 2: Test design: The test generator then devises a “malicious plan” by specifying the target negative consequence and the website to exploit. A key auxiliary consideration here is the time gap: the threat always occurs after the knowledge cutoff of the backend LLMs. This highlights the time-sensitive nature of the threat. To illustrate, testing whether agents will misbehave with a single unreliable and long-outdated website, alongside several current sources, arguably constitutes an unfaithful and unrealistic modeling of the threat. Notably, we do not bind a test case to a specific date during generation, ensuring that the resulting templates remain reusable rather than one-off applications.
@@
    Step 3: Test instantiation: The test generator turns each verbalized test proposal into (1) a set of guidelines for the website generator and (2) a checklist for safety evaluation. The guidelines specify the required format and the key facts the synthesized unreliable website should contain. In our implementation, we do not explicitly instruct the website to be persuasive. Inspired by rubric-based evaluation (Souly et al., 2024), the checklist enables precise assessment of whether the agent’s final response exhibits the targeted negative consequence.
@@
Automated quality filtering. The absence of unsafe responses may simply indicate an infeasible test case, rather than a resilient agent, creating a false sense of safety. To avoid this, we define two filtering criteria for generated test cases via differential testing with a baseline agent under two settings: a benign tool and a manipulated tool (see Section 3.3). The criteria are: (1) Attainability: Under the manipulated tool setting, we run a dry test with a baseline search agent to verify that the intended consequence is attainable. (2) Integrity: Under the benign tool setting, the agent should not trigger the target consequence. A test case that satisfies both criteria is realistic and actionable.
3.3 Simulation-Based Testing & Evaluation Methods
@@
Given a user query $u$, the agent $S$ issues a sequence of search tool calls $\{s_{1},s_{2},\dots,s_{n}\}$ to collect information and then synthesizes a final response. As illustrated in Figure 4, we simulate unreliable search results by injecting a single unreliable website $d_{u}$ into the returned list. When $n>1$, we inject $d_{u}$ only in the first search call, giving the agent an opportunity to down-weight or disregard it in subsequent calls. Thus, the final search results presented to $S$ are $D\cup\{d_{u}\}$, where $D=\{d_{1},d_{2},\dots,d_{k}\}$ denotes the authentic results. We generate $d_{u}$ from predefined guidelines and condition it on the testing date, mirroring the reality that search services continually surface up-to-date (and potentially unreliable) content. By mixing one unreliable website with authentic sources, our red-teaming framework approximates realistic threat scenarios while preserving the agent’s normal execution and response synthesis.
@@
We evaluate the agent by auditing its responses. Following prior studies in agent safety evaluation (Ruan et al., 2024; Zhang et al., 2024; Yin et al., 2024), we use an LLM-as-a-Judge (Zheng et al., 2023) for automated evaluation.
@@
    Safety evaluation. The evaluator receives the user query, target consequence (with the checklist), and agent response, engages in a reasoning process, and then delivers a final boolean judgment (see Appendix M). This yields a signal of potential compromise of the search agent. Across all trials, we derive a safety metric for search agents, which we term the Attack Success Rate (ASR). We validate checklist-assisted evaluation via human study in Appendix G, with human agreement of over 95%.
@@
    Helpfulness evaluation. We assess user-perceived helpfulness, i.e., how useful or plausible the agent’s response appears, rather than its factual accuracy. Given the user query and agent response, an evaluator assigns a score from 1 (not helpful) to 5 (extremely helpful) using Appendix M. We rescale scores to 0–100 and average across trials to obtain the Helpfulness Score (HS).
Figure 4: Illustration of red-teaming via simulation-based testing.
4 Experiments
In this section, we describe our experimental setup, present two key utilities of SafeSearch: benchmarking (Section 4.2) and agent auditing (Section 4.3), and demonstrate how it provides continuous supervision (Section 4.4).
4.1 Experimental Setup
Test case synthesis. Following Section 3.2, we curate 300 test cases spanning the five risk types, establishing the SafeSearch dataset. This is accomplished by iteratively running the generate–filter process until we obtain 60 test cases for each risk. We describe dataset characteristics, validate test case representativeness, and compare user queries to other benchmarks in Appendix C. We also ensure query integrity without unreliable websites in Appendix I.
Table 1: Main results of search agents on SafeSearch. Reasoning models are highlighted in light blue, while non-reasoning models are shown in light beige. We use bold to denote the best performance and underline for the second best. Timestamp: Sep 9, 2025. The experiments for Claude-4.5 models were conducted on Nov 15, 2025, with search results reused when available.
@@ 4.1 Experimental Setup
Search agents. We evaluate three representative search agent implementations with a comprehensive set of backend LLMs; details are in Section D.1.444As our red-teaming method involves hooking search tools, we exclude proprietary search agent systems. However, the method can naturally be adopted by internal testers of these systems.
@@
    LLM w/ search workflow. LLMs passively accept the search results of the verbatim user query before generation. This is adopted from FreshLLMs (Vu et al., 2024).
@@
    LLM w/ tool calling. The agent may proactively issue search queries as needed, with up to three tool calls.
@@
    Deep research scaffold. We build on the prototype from Google [^5] https://github.com/google-gemini/gemini-fullstack-langgraph-quickstart. This system includes query decomposition, one-turn search, reflection, and answer finalization, with a maximum of 3 research loops and 2 queries per round.
Models. Since search agents must handle long website contexts, we include 9 proprietary and 8 open-source LLMs with at least 32K token context windows and native tool-calling support. By default, we use the chat or instruction-tuned version and enable the reasoning mode if supported. See Section D.2 for model details.
Red-teaming configurations. Balancing cost and effectiveness, we employ o4-mini as the test generator, while GPT-4.1-mini serves as the website generator, safety evaluator, and helpfulness evaluator. For automated filtering, we use Qwen3-8B as the baseline model. The search tool returns the top five results with webpage content by default [^6] We use the Serper API (https://serper.dev/) for real-time Google Search results and the Jina Reader API (https://jina.ai/reader/) for LLM-friendly content extraction., and each website is limited to 2,000 tokens (measured with the GPT-5 tokenizer). The unreliable website is always appended to the result list. We cache search results across agents for fairness and to reduce stochasticity. The LLM temperature is set to 0.6, and each test case is run three times, with ASR and HS averaged across trials. All relevant prompts are provided in Appendix M.
Due to space limit, we leave the validation of seven testing choices to Appendix F: injection position, per-website token limit, website generation timestamp, website generator choice, injection round in multi-turn cases, advanced website generation strategies (e.g., enforcing persuasiveness), and choice of test generators.
@@ 4.2 Main Findings of Red-Teaming with SafeSearch
Search agents are easily misled by unreliable search results. Across varying settings, most search agents exhibit alarmingly high ASRs. This suggests that, once unreliable results enter the context, agents often fail to discount them and instead propagate risky content to users. Even the safest configuration, Claude-Sonnet-4.5 with tool-calling, still has a non-zero ASR of 4.6%. Qualitative analyses (Appendix L) indicate that the dominant failure mode is uncritical trust in search results.
Backend LLM choice strongly affects search agent safety. Search agent safety depends on how the backend LLM interprets unreliable search results. Under the same scaffold, different LLMs show substantially different resilience. Overall, reasoning models tend to be more robust than non-reasoning ones, as reflected by lower ASRs across categories. Notably, Gemini-2.5-based agents do not consistently outperform others despite their later knowledge cutoff. Interestingly, the relative rankings of backend LLMs are also largely stable across the search-workflow and tool-calling settings.
Agent scaffolds also play a critical role in safety. For instance, GPT-4.1-mini attains an overall ASR of 90.5% under the search workflow, which drops to 77.8% with tool-calling and further to 57.4% with the deep-research scaffold. This gap suggests that safety is not solely a model property. Long-horizon scaffolds (tool-calling or deep research) can mitigate the threat by enabling cross-validation across sources, issuing follow-up searches to resolve conflicts, and reducing reliance on any single (potentially unreliable) source. These benefits come with a cost trade-off: stronger scaffolds typically require more tool calls and longer contexts, which may increase latency and expense.
@@
Susceptibility varies substantially across risk types. On average, ASR differs markedly across risks. For example, GPT-5 with tool-calling achieves 0% ASR for Indirect Prompt Injection but remains vulnerable to other risks. Overall, Indirect Prompt Injection has a lower average ASR, indicating stronger resilience to this security threat (Wallace et al., 2024; Debenedetti et al., 2025). By contrast, Misinformation is the most severe threat.
@@
Figure 5: In-depth analyses of SafeSearch. Experiments are on LLMs with a search workflow. (Left) SafeSearch enables understanding of safety progress in agent development. Search result count correlates with agent susceptibility to unreliable results. Timestamp: Sep 17, 2025. (Middle) SafeSearch reveals defense effectiveness. Timestamp: Sep 17, 2025. (Right) The effectiveness of test cases is not dependent on the generation timestamp of the unreliable website.
Safety and helpfulness can be improved simultaneously with proper orchestration. Notably, unsafe behavior does not necessarily reduce perceived helpfulness. For instance, under tool calling, the average HSmanip. (92.2) exceeds HSbenign (91.4). This indicates that unsafe influence can be subtle, yielding responses that still appear useful to end users. Moreover, some agent designs achieve strong safety and helpfulness simultaneously, as illustrated by the GPT-5 series (see Appendix H). This effect is especially pronounced under cost scaling: the deep-research scaffold narrows capability gaps across LLMs, enabling even smaller models such as Qwen3-8B to remain both useful and safer.
4.3 SafeSearch Tracks Safety in Agent Development
Beyond benchmarking, SafeSearch provides a viable way to understand the implicit impacts of agent configurations on safety, while also uncovering the real effectiveness of defense strategies.
The impact of search result count on safety. Agent developers can configure how many search results are provided to LLMs, balancing utility against cost efficiency. As shown in Figure 5 (Left), our empirical findings demonstrate that this choice can implicitly influence the safety performance of search agents. When fewer search results are considered, unreliable websites have more chances to be included in the final response, thereby increasing the ASR. This analysis also serves as an ablation study of our red-teaming setup, showing that the threat is not limited to specific search result count and that a default choice of five is representative.
@@ 4.3 SafeSearch Tracks Safety in Agent Development
The effectiveness of defense strategies. We consider two baseline defenses: 1) Reminder, which augments the agent’s prompt with a warning about potentially unreliable websites; 2) Filtering, which applies an auxiliary detector (i.e., GPT-4.1-mini) to filter unreliable search results before returning them (Appendix M). As shown in Figure 5 (Middle), both defenses reduce risks to some extent but fail to eliminate them. The reminder prompting method, which reflects a common strategy of adding caveats in system prompts, proves largely ineffective. Filtering is more effective, reducing ASR by roughly half across two search agents, consistent with its proactive design and a recall of 44.2%. A retrospective manual review of the detector’s reasoning reveals that most detection successes arise from the relative lack of stealth in unreliable websites. To explain, such stealthiness is not currently a principle considered in our test case generation. In other words, specially crafted stealthier websites can pose greater threats, as studied in prior works (Chen and Shu, 2024). Overall, our red-teaming framework enables evaluating defense effectiveness and fosters greater transparency in search agent development.
4.4 SafeSearch Enables Continuous Supervision
In the preceding two sections, we primarily used 300 frozen test cases. Here, we highlight SafeSearch’ key advantage as a red-teaming framework: it can continuously supervise search agent safety in the long run, achieved by:
  • $\bullet$
    Curating more challenging test cases. We can achieve this by generating new test cases against increasingly stronger baseline agents. As shown in Appendix E, test cases that defeat stronger baseline agents (GPT-5-mini) achieve persistently higher ASRs than those that only defeat weaker agents (Qwen3-8B).
  • $\bullet$
    Configuring newer timestamps. A single test case, as a template, can be replicated into multiple distinct tests by varying the timestamp used to generate unreliable websites. As shown in Figure 5 (Right), the effectiveness of an unreliable website decreases only slightly, demonstrating that the test is not solely dependent on a specific generation timestamp and can serve for long-term use.
  • $\bullet$
    Efficacy under evolving authentic sources. As shown in Figure 9, even with a fixed timestamp and unreliable website, the test remains effective when paired with more up-to-date authentic websites.
5 Discussions
Simulation-based testing. Similar to Ruan et al. (2024); Xu et al. (2025a), we adopt a simulation-based method to evaluate search agent safety. This choice is arguably necessary, as arbitrarily manipulating real-world search engine rankings (Kumar et al., 2019) could harm other normal users and pose ethical risks. By contrast, simulation-based testing enables reproducible and cost-efficient evaluation without external side effects. We further discuss how simulation-based testing relates to real-world threats and outline our efforts to ensure representativeness in Section J.3.
Other safety layers. This work centers on agent-level safety. At the same time, detecting unreliable websites at the search service level is also a valuable and complementary mitigation strategy (see Section J.4 for details). Together, these approaches form multiple safety layers: search agents determine how search results are safely interpreted and acted upon, while search services contribute to safety by influencing which information is surfaced in the first place.